Cycle Time and Velocity

Norton Company has the following data for one of its production departments:

Theoretical velocity: 180 units per hour
Productive minutes available per year: 62,000,000
Annual conversion costs: $372,000,000
Actual velocity: 72 units per hour

Required:

1. Calculate the actual conversion cost per unit using actual cycle time and the standard cost per minute. Round your actual cycle time answer to three decimal places and your cost per unit answer to the nearest cent.

Actual cycle time minutes per unit
Standard cost per minute $ per minute
Conversion cost per unit $ per unit

2. Calculate the ideal conversion cost per unit using theoretical cycle time and the standard cost per minute. If required, round your intermediate calculations and final answers to two decimal places.

Theoretical cycle time minutes per unit
Conversion cost per unit $ per unit

What incentive exists for managers when cycle time costing is used?

  • To reduce cycle time
  • To increase cycle time
  • To keep cycle time constant

3. What if the actual velocity is 126 units per hour? What is the conversion cost per unit? If required, round your intermediate calculations and final answers to two decimal places.

Actual cycle time minutes per unit
Conversion cost per unit $ per unit
Answer #1

SOLUTION

1. Actual cycle time = 60 mins / 72

= 0.833 minutes per unit

(Cycle time is reciprocal to velocity)

Standard cost per minute = Annual conversion cost / Productive minutes available per year

= $372,000,000 / 62,000,000

= $6 per minute

Conversion cost per unit = Standard cost per minute *  Actual cycle time

= $6 * 0.833 = $5 per unit

2. Theoretical cycle time = 60 minutes / 180 = 0.33 minutes per unit

Conversion cost per unit = Standard cost per minute * Theoretical cycle time

= 6 * 0.33 = $1.98 per unit

- To reduce cycle time because it reduces the cost per unit

3. Actual cycle time = 60 mins / 126

= 0.48 minutes

Conversion cost per unit = $6 * 0.48

= $2.88 per unit